### v1

# FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models

FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models
Hongyang Wang1, Yichen Shi2, Zhuofu Tao3, Yuhao Gao1, Liepiao Zhang4, Xun Lin5
Jun Feng1, Xiaochen Yuan6, Zitong Yu5🖂, Xiaochun Cao7
1Shijiazhuang Tiedao University   2Shanghai Jiao Tong University  3UCLA  4GRGBanking
5Great Bay University   6Macao Polytechnic University   7Shenzhen Campus of Sun Yat-sen University
Equal contribution      🖂 Corresponding author
Figure 1: FaceShield Multi-task Response Demonstration. This figure illustrates the model’s performance across four tasks: coarse-grained classification, fine-grained classification, reasoning, and localization. Coarse-grained classification distinguishes real from spoofed faces, while fine-grained classification identifies specific attack types, such as print attacks. The reasoning task provides explanations for spoofing by analyzing features like lighting and symmetry. Localization precisely detects spoofed regions. This illustration highlights FaceShield’s capability to handle a range of complex questions with accuracy.
Face anti-spoofing (FAS) is crucial for protecting facial recognition systems from presentation attacks. Previous methods approached this task as a classification problem, lacking interpretability and reasoning behind the predicted results. Recently, multimodal large language models (MLLMs) have shown strong capabilities in perception, reasoning, and decision-making in visual tasks. However, there is currently no universal and comprehensive MLLM and dataset specifically designed for FAS task. To address this gap, we propose FaceShield, a MLLM for FAS, along with the corresponding pre-training and supervised fine-tuning (SFT) datasets, FaceShield-pre10K and FaceShield-sft45K. FaceShield is capable of determining the authenticity of faces, identifying types of spoofing attacks, providing reasoning for its judgments, and detecting attack areas. Specifically, we employ spoof-aware vision perception (SAVP) that incorporates both the original image and auxiliary information based on prior knowledge. We then use an prompt-guided vision token masking (PVTM) strategy to random mask vision tokens, thereby improving the model’s generalization ability. We conducted extensive experiments on three benchmark datasets, demonstrating that FaceShield significantly outperforms previous deep learning models and general MLLMs on four FAS tasks, i.e., coarse-grained classification, fine-grained classification, reasoning, and attack localization. Our instruction datasets, protocols, and codes will be released soon.
1 Introduction
Face anti-spoofing (FAS) is essential in facial recognition systems, ensuring that presentation attacks (PAs), such as print, replay, and 3D wearable masks, are effectively prevented. It has attracted considerable interest in industry and academia in the past decade.
Existing deep learning FAS models can be categorized into two types: vision-based methods and vision-language-based methods. As shown in Fig. 2(a), vision-based methods rely solely on image data (e.g., RGB, Depth, Infrared(IR)) and binary labels to train CNNs [66, 67] or ViTs [15, 70] for FAS. While they can achieve satisfactory results against known attack types and environments, these methods are prone to overfitting on spurious correlations and lack strong extrapolation capabilities. As illustrated in Fig. 2(b), Vision-language-based methods do not use binary labels but instead train CLIPs with image-text pairs [55, 36, 43]. The text labels in these methods provide more domain-agnostic information, enhancing models’ generalization capability. Although these FAS models demonstrate some recognition capabilities, they still face challenges such as limited generalization ability, poor interpretability, and a lack of capability for fine-grained localization of attack regions.
Recently, MLLMs have shown remarkable capabilities across various visual tasks, such as remote sensing [76, 29, 44], medical imaging [32, 57], and deepfake detection  [64, 25]. By leveraging the general capabilities of language foundation models alongside the visual information extracted by vision towers, these specialized MLLMs integrate perception, reasoning, and decision-making within a single model. Regarding the FAS task, SHIELD [53] conducted extensive evaluations on existing general-purpose MLLMs, revealing that their performance on FAS tasks still has room for improvement. [75] introduced a model capable of performing classification and description attack type. However, the model is limited in its ability to handle more nuanced tasks, such as identifying specific attack types, reasoning, and localizing spoofed areas. These limitations underscore the broader challenges in training FAS MLLMs, including: (1) a lack of pretraining and SFT datasets specific to FAS tasks, (2) the need to extend traditional FAS tasks to fully exploit MLLM capabilities, and (3) the difficulty for general-purpose vision towers to capture the subtle distinctions between real faces and PAs, unlike with natural images.
Motivated by the above discussion, in this paper, we expand the traditional FAS task to include four sub-tasks (see Fig. 1 for examples): coarse-grained classification, fine-grained classification, reasoning, and attack localization. We then introduce FaceShield, an MLLM specifically designed for these tasks. As can be seen from Fig. 2(c), we propose a pretraining and SFT dataset generation pipeline. This pipeline constructs two multimodal FAS instruction datasets containing 50k dialogues for FaceShield training. To the best of our knowledge, FaceShield is the first FAS MLLM, equipped with multiple detection capabilities. Additionally, FaceShield-pre10K and FaceShield-sft45K are the first high-quality datasets that can be used to train a FAS-specific MLLM. Our main contributions include:
  • We develop a novel data generation pipeline that utilizes a MLLM and predefined prompts, and construct two multimodal FAS instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K) with 12 attack types. To our best knowledge, these are the first multitask instruction datasets for the FAS community.
  • We propose FaceShield, the first multitask MLLM for FAS that is capable of coarse-grained classification, fine-grained classification, reasoning, and attack localization. FaceShield  utilizes the Spoof-Aware Vision Perception (SAVP) and Prompt-guided Visual Token Masking (PVTM) strategies to enhance the discrimination of confusing attack areas.
  • Extensive experiments demonstrate that FaceShield significantly outperforms previous specialized FAS models and general MLLMs across multiple datasets in various FAS evaluation tasks.
Figure 2: Pipelines of different FAS methods (a) traditional deep learning models, (b) multimodal models, and (c) MLLM
2 Related Work
Figure 3: Construction pipeline of our proposed instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K).The initial datasets (WMCA, PADISI, SiW-Mv2) are combined to form a uni-class dataset covering 12 spoofing types, with selected images annotated for visual grounding. Using MLLM with structured prompts, we generate two datasets: a pretraining dataset and an SFT dataset divided into four tasks (coarse-grained classification, fine-grained classification, reasoning, and localization). The pretraining data is filtered by CLIP for similarity, producing the FaceShield-pre10k dataset. SFT data undergoes multi-level filtering (LLM-based, keywords, and human reviews), followed by augmentation, resulting in the FaceShield-sft45k dataset.
Face Anti-Spoofing. Previous works usually employed convolutional neural networks (CNNs) [66, 67] or vision transformers (ViTs) [15, 70, 10] as backbone networks, integrating auxiliary priors for live/spoof classification (e.g., reflection light [6], depth reconstruction [6, 41, 3], rPPG signals [33, 65, 41, 73], and gradient information [68, 8]) to detect PAs. While these methods perform well in source domains, their performance degrades significantly in unseen domains due to domain shifts and unseen attacks [71]. Recently, many domain-generalized FAS methods [71] have been proposed to address these cross-domain challenges. Adversarial training [27, 26] and feature disentanglement techniques [42, 28, 17, 62] have been explored to decouple image features into domain-invariant (content) and domain-specific (style) features, preventing the model from learning spurious correlations between domain-specific features and labels. Some approaches employ meta-learning [48, 52, 7], data augmentation [9], or data synthesis methods [14, 39] to simulate domain shifts, enabling the model to learn a broader data distribution during training. Others perform domain [26, 58, 21, 60, 78] or gradient [30] alignment to constrain the relative relationships between sample embeddings across different source domains, aiming to learn more generalized decision boundaries. More recently, some methods have adopted vision-language models (VLMs), particularly CLIP [49], to detect PAs using text guidance. By leveraging the reasoning capabilities acquired during large-scale pretraining, VLM-based approaches [56, 36, 61, 37, 40] have shown to improve the generalization of FAS. Unlike traditional binary live/spoof classification, VLM-based methods classify images by computing similarities between visual features and textual descriptions of live and spoof faces, leveraging more causal information [36].
Existing FAS methods lack interpretable reasoning capabilities, e.g., language-based explanations, to distinguish between live and spoof faces. A concurrent work [75] collect 12 datasets and described images based on attack types and keywords. Using instruction tuning, they develop a MLLM capable of classifying and providing detailed descriptions of attacks. Meanwhile, these approaches fail to integrate classification, localization, and reasoning tasks simultaneously, limiting their adaptability and reliability in real-world applications, where versatile understanding is critical for reliable decision-making. Additionally, existing FAS datasets [61, 37] lack annotations for reasoning processes, further hindering the development of reasoning-capable FAS methods.
Figure 4: Comparison of different model architectures. (a) Single modal baseline model. (b) Proposed model without Prompt-Guided Vision Token Masking (PVTM). (c) Proposed model with PVTM. The input image is first encoded by the vision encoder, and PVTM partitions the encoded features into RGB, SAV, and Masked tokens. These visual tokens are passed to the projector, while the prompt is tokenized separately, producing text tokens. The combined visual and text tokens are then fed into the Large Language Model (LLM) with LoRA, producing the final output.
Multimodal Large Language Model. MLLMs, such as GPT-4V [2], LLaVA [38], Bunny [20] have demonstrated successful performances in general domains [13, 31, 24], attracting considerable attention from researchers. Researchers have developed task-specific MLLMs in various fields under vision tasks, including remote sensing [76, 29, 44], medical [32, 57], aesthetic evaluation [23], deepfake detection [64, 25] and Visual Grounding [4, 63, 46]. Unlike traditional models, MLLMs can handle multiple tasks within a single model and exhibit strong performance in unknown domains. FakeShield [64] proposes an explainable multi-task model for image forgery detection and localization, capable of making judgments based on tampering clues at both the image and pixel levels. The language foundation model contains scarce knowledge of face attacks, resulting in poor performance of existing MLLMs on FAS tasks. Our constructed datasets can introduce abundant FAS-related knowledge to MLLM, which through SAVP further enhances its ability to distinguish between real faces and PAs, and PVTM enhances the generalizability of the extracted features.
3 FaceShield-pre10K and FaceShield-sft45K
3.1 Dataset Collection
Fig. 3(a) shows the collection and localization coordinate annotation process of existing FAS datasets. Based on class types of WMCA (W) [16], PADSIS (P) [51] and SiW-Mv2 (S) [18] datasets, we organize all annotation categories into 12 distinct types: Bonafide, Fakehead, Print, Glasses, Replay, Paper mask, Flexible mask, Rigid mask, Partial Eye, Partial Mouth, Makeup, and Tattoo. We have also re-annotated the original dataset images at both the image and region levels. Finally, we obtained 12,091 images with class labels and 3,139 images with bounding box labels. More details and examples can be found in our Appendix.
3.2 Instruction Construction
After collecting images, class labels, and attack regions, as shown in Fig. 3(b), we use Bunny-Llama-3-8B-V [20] to assist in generating the pretrain dataset and the SFT dataset. Specifically, we first use a system prompt as a preamble to enhance the performance of responses from the MLLM assistant. Class labels are embedded in the prompt, providing descriptions of some image content. Subsequently, given the task type, the MLLM assistant generates multiple question-and-answer pairs for the task, following the format provided in the few-shot examples.
For the construction of the pretraining dataset FaceShield-pre10K, we only use the MLLM assistant to generate descriptions of the images, without any other information. The image and its corresponding generated description are filtered using a CLIP [50] score, removing image-text pairs with a similarity of less than 15%.
For the task-specific instruction dataset FaceShield-sft45K, after generating question-answer (QA) pairs using MLLM, manual filtering, and specific keyword filtering are applied to remove incorrect QA pairs. We enhance the diversity of the content in questions and answers using Llama3 [12] on the golden seed dataset obtained after filtering, to improve the dialogue capabilities of the model trained on this dataset. The specific content and responses for each task are as follows:
  • Coarse-grained classification: MLLM directly responds with labels for face images, identifying them as either real faces or PAs.
  • Fine-grained classification: The responses from the MLLM include not only real or PAs but also judgments on the specific categories of the PAs.
  • Reasoning: MLLM provides descriptions and reasoning from multiple perspectives on the image and delivers the final judgment.
  • Attack localization: MLLM first evaluates the faces in the image, if PA is detected, it provides the specific coordinates of the attack area.
4 FaceShield
Our goal is to train a FAS task-specific MLLM with two main objectives: 1) Enhance the visual encoder’s ability to extract features from real faces and presentation attacks, and 2) Utilize the extensive knowledge stored in the LLM to improve the model’s generalization capabilities when facing unknown domains. A naive training approach is shown in Fig.  4(a), where direct pre-training and SFT are conducted using RGB images and constructed QA data. However, the high similarity between real faces and PAs in RGB appearance poses significant challenges to this method. As shown in Fig. 4(b), Spoof-Aware Vision Perception (SAVP) combines images preprocessed based on prior knowledge, by extracting predefined local descriptor operators [72], with the original RGB image. Our complete model framework is shown in Fig. 4(c), RGB images and the extracted local descriptor images are fed into the vision encoder to extract vision token $V_{RGB}$ and $V_{SAV}$, respectively. These features are then processed through the Prompt-Guided Vision Token Masking (PVTM) module, which extracts highly generalizable vision tokens. These tokens are sent to a projector to align with text prompt token $P$ to produce $V_{align}$, which is then fed into the language foundation model for inference result $\mathcal{Y}$ as follows:
$V_{align}=\text{Projection}(V_{RGB},V_{SAV})$ (1)
$\mathcal{Y}=\text{MLLM}(V_{align},P)$ (2)
4.1 Spoof-Aware Vision Perception
As shown in Fig. 4(a), bonafide faces and PAs lack distinct discriminative features in RGB-based appearance space, whereas local descriptors [66, 72] extracted through image preprocessing can enhance their subtle live/spoof clues. As shown in Fig. 4(b), we extract features from the original images using Local Binary Pattern (LBP) [47], Gray, and Histogram of Oriented Gradients (HOG) [11], and concatenate them. LBP and Gray-specific computations are as follows:
$LBP$ $=\sum_{i=0}^{P-1}s(g_{i}-g_{c})\cdot 2^{i},$ (3)
$s(x)$ =$=\begin{cases}1&\text{if }x\geq 0,\\ 0&\text{otherwise},\end{cases}$
where $g_{c}$ is the pixel value of the central pixel in the considered neighborhood, and $g_{i}$ represents the pixel values of the $P$ surrounding pixels.
$\text{Gray}(I)=0.299\cdot R+0.587\cdot G+0.114\cdot B,$ (4)
where $R$, $G$, and $B$ are the red, green, and blue intensity values of the pixel, respectively.
The HOG calculates the gradient magnitude and direction at each pixel using edge detection operators and then divides the image into small, overlapping cells. Within each cell, gradients are binned according to their direction into histograms. Histograms from the cells within each block are concatenated and normalized based on the block’s overall gradient energy. The final HOG descriptor is formed by the vector of these normalized histograms from all blocks.
We perform the above three steps of feature extraction on the original image, then concatenate these as three channels to form a complete image. This composite image is then fed into the vision encoder to extract features as complementary information. The spoof-aware vision token $V_{SAV}$ and final vision input $V$ for FaceShield as follows:
$V_{SAV}=\text{Encoder}(\text{Concat}[LBP,Gray,HOG])$ (5)
$V=\text{Concat}[V_{RGB},V_{SAV}])$ (6)
Figure 5: Prompt-guided Vision Token Masking Module (PVTM)
4.2 Prompt-Guided Vision Token Masking
To further enhance the alignment between visual features and text prompts, and alleviate overfitting on spurious correlations, we leverage text prompts to guide token selection after the visual encoder. As shown in Fig. 5, the text tokens extracted from the prompt are pooled [54], and then their similarity with all visual tokens is calculated. We assume that visual tokens with higher similarity are more relevant to subsequent tasks and even for the same image, the important visual tokens may not be consistent across different tasks. The calculation of similarity between visual tokens $V_{i}$ and text prompt tokens $P$ is as follows:
$Sim(V_{i},P)=\frac{V_{i}\cdot P}{\|V_{i}\|\|P\|}$ (7)
Subsequently, we apply a softmax function to the similarities between all $V_{i}$ and $P$, using the resulting values $S^{i}_{rank}$ as an importance metric for each visual token. We then rank all tokens based on this importance calculated as follows:
$S^{i}_{rank}(V_{i},P)=\frac{e^{S(V_{i},P)}}{\sum_{j}e^{S(V_{j},P)}}$ (8)
Afterward, we retain the top $k\%$ of vision tokens in importance. The remaining tokens are then randomly masked with a probability of $p\%$, reducing the influence of less important tokens while keeping acceptable information loss for the final decision-making process.
4.3 Training Details
We employ a two-stage training strategy. The first stage is the pre-training stage, where we perform continual pretrain of existing weights. In this stage, we align the visual embeddings from a pre-trained vision encoder with the text embeddings using FaceShield-pre10K. The second stage is the supervised fine-tuning (SFT) stage, where we apply visual instruction tuning to fully leverage the MLLM’s capabilities across various multimodal tasks within our specific domain using FaceShield-sft45K. For LLM tuning, we incorporate LoRA [22] and set the parameters with a hidden dimension of 128 and a scaling factor of 256 in every transformer block. In both stages, we employ cross-entropy loss for next-token prediction. During the pre-training stage, only the projector after the vision encoder and PVTM are optimized for one epoch. In the SFT stage, we fine-tune the LoRA layers in LLM as well as the projector.
Table 1: Intra-dataset results on coarse-grained classification.
Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
Traditional
ResNet [19] 97.55 2.32
PatchNet [59] 98.22 1.78
CoOp [77] 98.73 1.27
MLLM
LLaVA [38] 65.54 27.76
Qwen-VL [4] 51.94 38.70
Minigpt4 [79] 26.86 65.50
Bunny [20] 81.20 17.87
Bunny (fine-tuned) [20] 98.23 1.52
FaceShield (Ours) 99.41 ± 0.06 0.53 ± 0.06
5 Experiments and Results
5.1 Protocols and Evaluation Metrics
We select 10% of each source dataset in FaceShield-sft45K to form three test subsets (W, S, P). For the coarse-grained classification and fine-grained classificationtask, we conduct both intra-dataset and cross-dataset testing. For the reasoning, and attack localization tasks, we employ intra-dataset evaluation. The intra-dataset test uses all data from the source datasets for training and evaluation on the combined test sets (W&S&P). For cross-dataset testing, we select two datasets for pre-training and SFT, and then test using the test set from the remaining dataset. For example, we might train on W and S, and test using the test set from P. For the coarse-grained classification, fine-grained classification, and reasoning tasks, we use Half Total Error Rate (HTER) [69] or accuracy (ACC) as evaluation metrics. Additionally, we evaluate the reasoning process for the reasoning task using BLEU [45], ROUGE-L [34], and METEOR [5], and evaluate the attack localizationtask using AP@40 and AP@50.
5.2 Implementation Details
We use Siglip [74] as the visual encoder and Phi-3 [1] as the language foundation model. PVTM retains the top 10% of the most important tokens and randomly masks 5% of the tokens in the remaining 90%. Adam optimizer is used in the pretrain stage with a learning rate of $5\times 10^{-4}$. As for SFT stage, we decrease the learning rate to $2\times 10^{-4}$. All experiments are conducted on a single NVIDIA A100 GPU. Each experiment is repeated 10 times on the model, and the final results are reported as the mean ± standard deviation.
Table 2: Cross-dataset results on coarse-grained classification. $W$, $S$, and $P$ denote WMCA, SiW-Mv2, and PADISI, respectively.
Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
W & S P
ResNet [19] 46.12 50.00
PatchNet [59] 77.18 22.87
IADG [78] 72.96 27.01
FaceShield (Ours) 93.17 ± 0.22 6.37 ± 0.21
W & P S
ResNet [19] 53.36 49.16
PatchNet [59] 56.16 45.37
IADG [78] 57.20 42.81
FaceShield (Ours) 89.93 ± 0.15 10.3 ± 0.14
S & P W
ResNet [19] 74.01 29.75
PatchNet [59] 78.15 41.50
IADG [78] 78.55 26.27
FaceShield (Ours) 92.56 ± 0.08 5.71 ± 0.08
Table 3: Results of fine-grained classification task.
Method ACC(%) $\uparrow$
LLaVA [38] 16.39
Qwen-VL [4] 16.55
Minigpt4 [79] 19.51
Bunny [20] 27.03
Bunny (Fine-tuned) [20] 94.43
FaceShield (Ours) 95.81 ± 0.11
Table 4: Results of the reasoning task with metrics BLEU, ROUGE-L, METEOR, ACC, and HTER.
Method BLEU-1 (%) $\uparrow$ BLEU-2 (%) $\uparrow$ BLEU-3 (%) $\uparrow$ BLEU-4 (%) $\uparrow$ ROUGE-L (%) $\uparrow$ METEOR (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
LLaVA [38] 45.05 31.75 23.42 17.80 30.51 25.52 37.84 50.11
Minigpt4 [79] 17.85 7.85 3.53 1.94 27.54 21.88 33.86 50.00
Qwen-VL [4] 20.92 14.53 11.01 8.77 21.45 12.49 47.64 41.49
Bunny [20] 33.64 27.12 22.65 19.33 36.74 19.12 50.68 39.73
Bunny(fine-tuned) [20] 89.57 86.96 84.91 81.29 80.15 51.64 98.56 1.16
FaceShield (Ours) 90.89 ± 0.14 88.02 ± 0.15 85.75 ± 0.17 83.98 ± 0.19 82.98 ± 0.20 53.10 ± 0.16 99.29 ± 0.04 0.57 ± 0.04
5.3 Comparison with Existing Methods
5.3.1 Coarse-Grained Classification Task
For the coarse-grained classification task, we compare FaceShield with state-of-the-art FAS methods  [19, 59, 77, 78] and open-source MLLM models [38, 4, 79, 20].
Intra-dataset Testing. Table 1 demonstrates that Our FaceShield significantly outperforms three representative traditional FAS methods [19, 59, 77] in within-domain experiments. Moreover, our performance greatly exceeds the zero-shot capabilities of general MLLM. We also fine-tune the open-source MLLM (i.e., Bunny [20]), selecting RGB images and language data from the dataset to conduct experiments on Bunny. We find that FaceShield also surpasses the well-tuned MLLM (Bunny) with 1% HTER decrease.
Cross-dataset Testing. Table 2 shows the performance of FaceShield in cross-domain scenarios, where we trained on two out of three selected datasets and tested on one. FaceShield demonstrates performance far exceeding traditional FAS models in cross-domain scenarios. Under the S&PW protocol, it achieves the HTER of 5.72%, showcasing FaceShield’s strong generalization capabilities compared to traditional methods.
5.3.2 Fine-Grained Classification Task
Table 3 shows the results under the fine-grained classification task. For open-source MLLMs [38, 4, 79, 20], we incorporated 12 types of attacks into the prompt, allowing it to respond with the correct type. For the fine-tuned MLLM and our FaceShield, we selected keywords from the responses for evaluation. It is evident that supervised fine-tuning can significantly improve the model’s performance, with FaceShieldachieving the best results. Cross-dataset results can be found in the Appendix due to space constraints.
5.3.3 Reasoning Task
We also explore the models’ reasoning capacity and Table 4 displays the results for the reasoning task. General MLLMs perform poorly in both classification and reasoning. FaceShield significantly outperforms general MLLMs in both the reasoning process and judgment, and also exceeds open-source MLLMs (e.g., Bunny) fine-tuned on our dataset. FaceShield not only provides accurate results but also delivers detailed and correct reasoning, effectively enhancing the explainability of FAS methods.
5.3.4 Attack Localization Task
It is vital to locate the spoof regions for explainable FAS, and Table 5 presents the results for the attack localization task. Due to the scarcity of attack localization annotations in general pre-trained datasets, general MLLMs perform poorly on this task. In contrast, FaceShield shows excellent results, achieving over 95% for both AP@40 and AP@50. It accurately locates attack areas, providing new insights into attack region detection for FAS tasks.
Table 5: Results of the attack localization task with metrics AP@40 and AP@50.
Method AP@40 (%) $\uparrow$ AP@50 (%) $\uparrow$
Qwen-VL [4] 2.07 1.49
Lenna [63] 37.77 35.41
Sphinx [35] 47.86 46.30
Bunny [20] 73.50 71.65
Bunny(fine-tuned) [20] 92.30 89.71
FaceShield (Ours) 97.78 ± 0.21 95.60 ± 0.19
Table 6: Ablation results across different tasks.
SAVP PVTM Coarse-grained classification Fine-grained classification Reasoning Attack localization
ACC(%) $\uparrow$ HTER(%) $\downarrow$ ACC(%) $\uparrow$ ACC(%) $\uparrow$ HTER(%) $\downarrow$ AP@40(%) $\uparrow$ AP@50(%) $\uparrow$
$\times$ $\times$ 98.23 1.52 94.43 98.56 1.16 92.30 89.71
$\checkmark$ $\times$ 98.73 1.06 94.59 99.41 0.48 97.09 95.21
$\checkmark$ $\checkmark$ $\mathbf{99.41\pm 0.06}$ $\mathbf{0.53\pm 0.06}$ $\mathbf{95.81\pm 0.11}$ $99.29\pm 0.04$ $0.57\pm 0.04$ $\mathbf{97.78\pm 0.21}$ $\mathbf{95.6\pm 0.19}$
Figure 6: Comparison of performance after fine-tuning using our proposed dataset on LLaVA [38] and Bunny [20] models.
5.4 Ablation Study
Effectiveness of the proposed datasets. We conduct pre-training and SFT with our proposed FaceShield-pre10K and FaceShield-sft45K on LLaVA [38] and Bunny [20] to evaluate the efficacy and generalization of our instruction tuning datasets. The results in Fig. 6 demonstrate that fine-tuning on our dataset significantly enhances performance across three tasks for both LLaVA and Bunny. This demonstrates that our dataset can effectively enrich general MLLMs with substantial FAS-related knowledge and boost their capabilities in FAS tasks. It also validates the efficacy of our advanced dataset construction pipeline.
Figure 7: Ablation of visual token masking ratio $p$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
Effectiveness of SAVP. Results from the first two rows in Table 6 show that leveraging local descriptors as complementary visual inputs significantly improves performance across four tasks, particularly in the attack localization task, where AP@40 and AP@50 increased by 5.6% and 5.94%, respectively. It indicates that prior knowledge-based auxiliary information significantly enhances the model’s ability to distinguish easily confusable facial images. The local live/spoof details within the auxiliary data proves especially valuable for fine-grained attack region detection tasks.
Figure 8: Ablation of visual token retain ratio $k$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
Effectiveness of PVTM. It can be seen from the last two rows of Table 6 that the proposed PVTM provides reasonable improvements for FaceShield across three (coarse- and fine-grained classification, and attack localization) tasks. It indicates that masking less important tokens helps prevent the model from task-unrelated noises and spurious correlations. However, PVTM leads to a slight performance decrease on the reasoning task. This might be because masking partial visual tokens may compromise overall image perception and lose information for reasoning.
We further conduct studies of PVTM with different visual token masking ratios $p$, as shown in Fig. 7(a) and (b). Through our experiments for three (coarse- and fine-grained classification, and attack localization) tasks, we find that masking 5% of the tokens strikes an optimal balance between reducing spurious correlations and preserving essential information. However, we cannot find any improvement when masking more tokens due to severe information loss. Besides, we also investigate varying proportions of visual token retain ratio $k$, as illustrated in Fig. 8(a) and (b). We preserve 0%, 10%, 20%, and 30% of the tokens, respectively, and then randomly mask $p$=5the remaining tokens. The results show that retaining $k$=10visual tokens with strong importance achieves optimal performance. However, the more tokens are preserved, the poorer the model performs, suggesting that as the importance of tokens decreases, the likelihood of spurious correlations increases.
Figure 9: Importance visualization of (a) RGB vision tokens for classification tasks and (b) SAV tokens for attack localization task.
6 Visualization and Analysis
Here, we visualize the importance of RGB and SAV tokens. Fig. 9(a) displays the significance of visual tokens in relation to the prompt text for images from three different categories. For the coarse- and fine-grained classification tasks, it is evident that the tokens sensitive to prompt are mainly clustered in the facial regions. The features of these areas play a decisive role in determining the final outcome. Fig. 9(b) demonstrates how SAV tokens assist in the attack localization task. For this task, RGB tokens sometimes fail to provide accurate localization information. As shown in the left of Fig. 9(b), for the eye mask region, the importance visualization of RGB tokens is quite dispersed. However, in the SAV token visualization, the significance of the eye attack region is clearly highlighted. The right part of Fig. 9(b) shows how RGB and SAV tokens together provide supportive information for the attack localization task, including the eye mask attack area and the holding hands.
7 Conclusion
In this paper, we expand the FAS task into four sub-tasks: coarse-grained classification, fine-grained classification, reasoning, and attack localization, and propose the FaceShield, a specialized MLLM tailored for these FAS tasks. Considering the specific training data requirements of MLLM, we establish a pipeline for constructing datasets tailored for FAS task pre-training and supervised fine-tuning, resulting in the creation of the FaceShield-pre10K and FaceShield-sft45K datasets. This paper represents a preliminary exploration of MLLM for FAS task, and future works will focus on incorporating multiple visual modalities and refining the granularity of attack region localization.

### v2

@@
 FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models
-FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models
-Hongyang Wang1, Yichen Shi2, Zhuofu Tao3, Yuhao Gao1, Liepiao Zhang4, Xun Lin5
-Jun Feng1, Xiaochen Yuan6, Zitong Yu5🖂, Xiaochun Cao7
-1Shijiazhuang Tiedao University   2Shanghai Jiao Tong University  3UCLA  4GRGBanking
-5Great Bay University   6Macao Polytechnic University   7Shenzhen Campus of Sun Yat-sen University
-Equal contribution      🖂 Corresponding author
-Figure 1: FaceShield Multi-task Response Demonstration.
-This figure illustrates the model’s performance across four tasks: coarse-grained classification, fine-grained classification, reasoning, and localization.
-Coarse-grained classification distinguishes real from spoofed faces, while fine-grained classification identifies specific attack types, such as print attacks.
-The reasoning task provides explanations for spoofing by analyzing features like lighting and symmetry.
-Localization precisely detects spoofed regions.
-This illustration highlights FaceShield’s capability to handle a range of complex questions with accuracy.
+Abstract
 Face anti-spoofing (FAS) is crucial for protecting facial recognition systems from presentation attacks.
@@
 We conducted extensive experiments on three benchmark datasets, demonstrating that FaceShield significantly outperforms previous deep learning models and general MLLMs on four FAS tasks, i.e., coarse-grained classification, fine-grained classification, reasoning, and attack localization.
-Our instruction datasets, protocols, and codes will be released soon.
-1 Introduction
+Our instruction datasets, protocols, and codes will be released at https://github.com/Why0912/FaceShield.
+Figure 1: FaceShield Multi-task Response Demonstration.
+This figure shows the model’s performance on four tasks: coarse-grained classification (real vs.
+spoofed faces), fine-grained classification (specific attack types like print attacks), reasoning (explaining spoofing using features such as lighting and symmetry), and localization (detecting spoofed regions).
+It highlights FaceShield’s ability to handle diverse, complex questions accurately.
+Introduction
 Face anti-spoofing (FAS) is essential in facial recognition systems, ensuring that presentation attacks (PAs), such as print, replay, and 3D wearable masks, are effectively prevented.
@@
 As shown in Fig.
-2(a), vision-based methods rely solely on image data (e.g., RGB, Depth, Infrared(IR)) and binary labels to train CNNs [66, 67] or ViTs [15, 70] for FAS.
+2(a), vision-based methods rely solely on image data (e.g., RGB, Depth, Infrared(IR)) and binary labels to train CNNs (yu2020searching; yu2021dual; wang2025pnss) or ViTs (9484333; yu2023rethinkingvisiontransformermasked; cai2025rehearsal) for FAS.
 While they can achieve satisfactory results against known attack types and environments, these methods are prone to overfitting on spurious correlations and lack strong extrapolation capabilities.
 As illustrated in Fig.
-2(b), Vision-language-based methods do not use binary labels but instead train CLIPs with image-text pairs [55, 36, 43].
+2(b), Vision-language-based methods do not use binary labels but instead train CLIPs with image-text pairs (Srivatsan_2023_ICCV; liu2024fmclip; mu2024tegdgtextuallyguideddomain; lin2025reliable).
 The text labels in these methods provide more domain-agnostic information, enhancing models’ generalization capability.
 Although these FAS models demonstrate some recognition capabilities, they still face challenges such as limited generalization ability, poor interpretability, and a lack of capability for fine-grained localization of attack regions.
-Recently, MLLMs have shown remarkable capabilities across various visual tasks, such as remote sensing [76, 29, 44], medical imaging [32, 57], and deepfake detection  [64, 25].
+Recently, MLLMs have shown remarkable capabilities across various visual tasks (ye2025cat), such as remote sensing (zhang2024earthgpt; kuckreja2023geochat; 2402.02544), medical imaging (li2023llavamed; Sun2024STLLaVAMedSL), and deepfake detection  (xu2024fakeshieldexplainableimageforgery; huang2024ffaa).
 By leveraging the general capabilities of language foundation models alongside the visual information extracted by vision towers, these specialized MLLMs integrate perception, reasoning, and decision-making within a single model.
-Regarding the FAS task, SHIELD [53] conducted extensive evaluations on existing general-purpose MLLMs, revealing that their performance on FAS tasks still has room for improvement.
-[75] introduced a model capable of performing classification and description attack type.
+Regarding the FAS task, SHIELD (shi2025shield) conducted extensive evaluations on existing general-purpose MLLMs, revealing that their performance on FAS tasks still has room for improvement.
+(zhang2025interpretable) introduced a model capable of performing classification and description attack type.
 However, the model is limited in its ability to handle more nuanced tasks, such as identifying specific attack types, reasoning, and localizing spoofed areas.
@@
 Figure 2: Pipelines of different FAS methods (a) traditional deep learning models, (b) multimodal models, and (c) MLLM
-2 Related Work
 Figure 3: Construction pipeline of our proposed instruction datasets (i.e., FaceShield-pre10K and FaceShield-sft45K).The initial datasets (WMCA, PADISI, SiW-Mv2) are combined to form a uni-class dataset covering 12 spoofing types, with selected images annotated for visual grounding.
@@
 SFT data undergoes multi-level filtering (LLM-based, keywords, and human reviews), followed by augmentation, resulting in the FaceShield-sft45k dataset.
+Additional details can be found in the appendix.
+Related Work
 Face Anti-Spoofing.
-Previous works usually employed convolutional neural networks (CNNs) [66, 67] or vision transformers (ViTs) [15, 70, 10] as backbone networks, integrating auxiliary priors for live/spoof classification (e.g., reflection light [6], depth reconstruction [6, 41, 3], rPPG signals [33, 65, 41, 73], and gradient information [68, 8]) to detect PAs.
-While these methods perform well in source domains, their performance degrades significantly in unseen domains due to domain shifts and unseen attacks [71].
-Recently, many domain-generalized FAS methods [71] have been proposed to address these cross-domain challenges.
-Adversarial training [27, 26] and feature disentanglement techniques [42, 28, 17, 62] have been explored to decouple image features into domain-invariant (content) and domain-specific (style) features, preventing the model from learning spurious correlations between domain-specific features and labels.
-Some approaches employ meta-learning [48, 52, 7], data augmentation [9], or data synthesis methods [14, 39] to simulate domain shifts, enabling the model to learn a broader data distribution during training.
-Others perform domain [26, 58, 21, 60, 78] or gradient [30] alignment to constrain the relative relationships between sample embeddings across different source domains, aiming to learn more generalized decision boundaries.
-More recently, some methods have adopted vision-language models (VLMs), particularly CLIP [49], to detect PAs using text guidance.
-By leveraging the reasoning capabilities acquired during large-scale pretraining, VLM-based approaches [56, 36, 61, 37, 40] have shown to improve the generalization of FAS.
-Unlike traditional binary live/spoof classification, VLM-based methods classify images by computing similarities between visual features and textual descriptions of live and spoof faces, leveraging more causal information [36].
-Existing FAS methods lack interpretable reasoning capabilities, e.g., language-based explanations, to distinguish between live and spoof faces.
-A concurrent work [75] collect 12 datasets and described images based on attack types and keywords.
-Using instruction tuning, they develop a MLLM capable of classifying and providing detailed descriptions of attacks.
-Meanwhile, these approaches fail to integrate classification, localization, and reasoning tasks simultaneously, limiting their adaptability and reliability in real-world applications, where versatile understanding is critical for reliable decision-making.
-Additionally, existing FAS datasets [61, 37] lack annotations for reasoning processes, further hindering the development of reasoning-capable FAS methods.
-Figure 4: Comparison of different model architectures.
-(a) Single modal baseline model.
-(b) Proposed model without Prompt-Guided Vision Token Masking (PVTM).
-(c) Proposed model with PVTM.
-The input image is first encoded by the vision encoder, and PVTM partitions the encoded features into RGB, SAV, and Masked tokens.
-These visual tokens are passed to the projector, while the prompt is tokenized separately, producing text tokens.
-The combined visual and text tokens are then fed into the Large Language Model (LLM) with LoRA, producing the final output.
-Multimodal Large Language Model.
-MLLMs, such as GPT-4V [2], LLaVA [38], Bunny [20] have demonstrated successful performances in general domains [13, 31, 24], attracting considerable attention from researchers.
-Researchers have developed task-specific MLLMs in various fields under vision tasks, including remote sensing [76, 29, 44], medical [32, 57], aesthetic evaluation [23], deepfake detection [64, 25] and Visual Grounding [4, 63, 46].
-Unlike traditional models, MLLMs can handle multiple tasks within a single model and exhibit strong performance in unknown domains.
-FakeShield [64] proposes an explainable multi-task model for image forgery detection and localization, capable of making judgments based on tampering clues at both the image and pixel levels.
-The language foundation model contains scarce knowledge of face attacks, resulting in poor performance of existing MLLMs on FAS tasks.
-Our constructed datasets can introduce abundant FAS-related knowledge to MLLM, which through SAVP further enhances its ability to distinguish between real faces and PAs, and PVTM enhances the generalizability of the extracted features.
-3 FaceShield-pre10K and FaceShield-sft45K
-3.1 Dataset Collection
+Early FAS methods primarily used CNNs (yu2020searching) and ViTs (9484333), incorporating auxiliary cues such as reflection (icassp22multiple), depth (cvpr18aux), and rPPG (yu2019remote) for live/spoof classification.
+While some fuse multi-modal inputs (e.g., RGB, IR, depth)(yu2024rethinking), they often suffer performance drops in unseen domains due to domain shifts(tpami23survey).
+To enhance generalization, domain-generalized FAS methods (tpami23survey) employ techniques such as adversarial training (ijcv23adversarial), feature disentanglement (cvpr22ssan), one-class learning (cvpr24oneclass), meta-learning (qin2021meta), data augmentation (ijcv24augmentation), and data synthesis (aaai22generation) to model domain shifts and separate domain-invariant content from domain-specific styles.
+Additionally, domain or gradient alignment (cvpr24gradient) aids in learning generalized decision boundaries for more robust and domain-invariant spoof detection.
+Recently, vision-language models (VLMs), especially CLIP (clip), have been explored for FAS (flip; liu2024fmclip).
+These models use textual descriptions of live and spoof faces to guide classification, offering stronger generalization by leveraging the powerful visual representations learned during pre-training (liu2024fmclip).
+However, most FAS methods lack interpretability, particularly in providing language-based explanations.
+A recent work (zhang2025interpretable) collected 12 datasets and used instruction tuning to train an MLLM for attack classification, but it struggles to integrate classification, localization, and reasoning tasks for real-world deployment.
+Additionally, existing FAS datasets (eccv25tffas; cvpr24cfpl) lack reasoning annotations, hindering the development of more comprehensive and interpretable models.
+Multimodal Large Language Models.
+MLLMs such as GPT-4V (gpt4), LLaVA (liu2023llava), and Bunny (he2024bunny) have shown strong performance across general domains (li2023seedbench2benchmarkingmultimodallarge; huang2024aesbenchexpertbenchmarkmultimodal).
+Task-specific MLLMs have also been developed for various vision tasks, including remote sensing (zhang2024earthgpt; kuckreja2023geochat), medical imaging (li2023llavamed; Sun2024STLLaVAMedSL), aesthetic assessment (huang2024aesexpert), deepfake detection (xu2024fakeshieldexplainableimageforgery; huang2024ffaa), and visual grounding (wei2023lenna; peng2023kosmos).
+Unlike traditional models, MLLMs unify multiple tasks within a single framework and generalize well to unseen domains.
+For example, FakeShield (xu2024fakeshieldexplainableimageforgery) introduces an explainable multi-task model for image forgery detection and localization by leveraging visual-textual clues at both global and pixel levels.
+However, current MLLMs lack specific knowledge about face attacks, leading to limited performance on FAS tasks.
+To address this, we introduce dedicated FAS datasets that enrich MLLMs with domain-specific knowledge.
+The proposed SAVP module enhances spoof discrimination by guiding the model with semantic attack priors, while PVTM improves feature generalization across varied domains.
+FaceShield-pre10K and FaceShield-sft45K
+Dataset Collection
+Figure 4: Proposed model architectures.
+(a) Proposed model with Spoof-Aware Vision Perception (SAVP).
+(b) Proposed model with SAVP and Prompt-Guided Vision Token Masking (PVTM).
+(c) Details about PVTM.
 Fig.
-3(a) shows the collection and localization coordinate annotation process of existing FAS datasets.
-Based on class types of WMCA (W) [16], PADSIS (P) [51] and SiW-Mv2 (S) [18] datasets, we organize all annotation categories into 12 distinct types: Bonafide, Fakehead, Print, Glasses, Replay, Paper mask, Flexible mask, Rigid mask, Partial Eye, Partial Mouth, Makeup, and Tattoo.
-We have also re-annotated the original dataset images at both the image and region levels.
-Finally, we obtained 12,091 images with class labels and 3,139 images with bounding box labels.
-More details and examples can be found in our Appendix.
-3.2 Instruction Construction
-After collecting images, class labels, and attack regions, as shown in Fig.
-3(b), we use Bunny-Llama-3-8B-V [20] to assist in generating the pretrain dataset and the SFT dataset.
-Specifically, we first use a system prompt as a preamble to enhance the performance of responses from the MLLM assistant.
-Class labels are embedded in the prompt, providing descriptions of some image content.
-Subsequently, given the task type, the MLLM assistant generates multiple question-and-answer pairs for the task, following the format provided in the few-shot examples.
-For the construction of the pretraining dataset FaceShield-pre10K, we only use the MLLM assistant to generate descriptions of the images, without any other information.
-The image and its corresponding generated description are filtered using a CLIP [50] score, removing image-text pairs with a similarity of less than 15%.
-For the task-specific instruction dataset FaceShield-sft45K, after generating question-answer (QA) pairs using MLLM, manual filtering, and specific keyword filtering are applied to remove incorrect QA pairs.
-We enhance the diversity of the content in questions and answers using Llama3 [12] on the golden seed dataset obtained after filtering, to improve the dialogue capabilities of the model trained on this dataset.
-The specific content and responses for each task are as follows:
-  • Coarse-grained classification: MLLM directly responds with labels for face images, identifying them as either real faces or PAs.
-  • Fine-grained classification: The responses from the MLLM include not only real or PAs but also judgments on the specific categories of the PAs.
-  • Reasoning: MLLM provides descriptions and reasoning from multiple perspectives on the image and delivers the final judgment.
-  • Attack localization: MLLM first evaluates the faces in the image, if PA is detected, it provides the specific coordinates of the attack area.
-4 FaceShield
-Our goal is to train a FAS task-specific MLLM with two main objectives: 1) Enhance the visual encoder’s ability to extract features from real faces and presentation attacks, and 2) Utilize the extensive knowledge stored in the LLM to improve the model’s generalization capabilities when facing unknown domains.
-A naive training approach is shown in Fig.
-4(a), where direct pre-training and SFT are conducted using RGB images and constructed QA data.
+3(a) illustrates the annotation process for existing FAS datasets.
+Based on the class types from WMCA (W) (8714076), PADSIS (P) (Rostami_2021_ICCV), and SiW-Mv2 (S) (xiaoguo2022MDFAS), we unify the annotation categories into 12 types: Bonafide, Fakehead, Print, Glasses, Replay, Paper mask, Flexible mask, Rigid mask, Partial Eye, Partial Mouth, Makeup, and Tattoo.
+We re-annotate all images at both image- and region-levels, resulting in 12,091 images with class labels and 3,139 images with bounding box annotations.
+The detail information is shown in Table 1.
+Instruction Construction
+As shown in Fig.
+3(b), we construct two instruction datasets using Bunny-Llama-3-8B-V (he2024bunny).
+A system prompt containing class labels is used to guide the MLLM assistant in generating question-answer (QA) pairs, based on the task type and few-shot examples.
+For the pretraining dataset FaceShield-pre10K, we generate image descriptions only, without task instructions.
+Pairs with a CLIP (radford2021learning) similarity score below 15% are filtered out to ensure quality.
+Algorithm 1 Construction of FaceShield-sft45K Dataset
+1:Image dataset $\mathcal{D}_{img}$ from WMCA, SiW-Mv2, PADISI;
+2:Multimodal LLM (MLLM); Task set $\mathcal{T}$.
+3:Processed QA dataset $\mathcal{D}_{QA}$.
+4:Initialize $\mathcal{D}_{QA}\leftarrow\emptyset$
+5:for image $I\in\mathcal{D}_{img}$ do
+6:  for task $t\in\mathcal{T}$ do
+7:   Generate task-specific prompt $\mathcal{P}_{t}$ using Ground Truth (GT), requirements and examples
+8:   Generate QA pair $(q,a)$ using MLLM: $(q,a)=\text{MLLM}(\mathcal{P}_{t},I)$
+9:   if QA pair $(q,a)$ satisfies validation criteria then
+10:     Add $(q,a)$ to $\mathcal{D}_{QA}$
+11:   end if
+12:  end for
+13:end for
+14:Filter low-quality QA pairs in $\mathcal{D}_{QA}$:
+15:  Remove pairs with low semantic consistency or invalid content
+16:  Perform human review to refine remaining pairs
+17:Augment $\mathcal{D}_{QA}$ by generating rephrased QA pairs:
+18:for QA pair $(q,a)\in\mathcal{D}_{QA}$ do
+19:  Generate additional pairs $(q^{\prime},a^{\prime})$ using rephrasing strategies
+20:  Add $(q^{\prime},a^{\prime})$ to $\mathcal{D}_{QA}$
+21:end for
+22:return $\mathcal{D}_{QA}$
+For the instruction-tuning dataset FaceShield-sft45K, as shown in Alg.
+1, MLLM-generated QA pairs are filtered using both manual and keyword-based strategies.
+The resulting high-quality seed set is then diversified using LLaMA3 (dubey2024llama) to enhance linguistic richness and dialogue ability.
+Table 1: Datasets and QA statistics
+Datasets Attack Types Annotations QA_Count
+FaceShield-pre10K - Image-text pairs 9297
+FaceShield-sft45K Unified-attack(11 types) QA + bbox 45662
+The dataset covers four tasks: (1) Coarse-grained classification, which predicts whether a face is real or spoofed; (2) Fine-grained classification, which identifies specific PA types beyond binary classification; (3) Reasoning, which provides explanations and justifications before making a judgment; (4) Attack localization, which outputs coordinates of attack regions if spoofing is detected.
+FaceShield
+Our goal is to train a FAS task-specific MLLM with two main objectives: 1) Enhance the visual encoder’s ability to extract features from real faces and presentation attacks, and 2) Utilize the extensive knowledge stored in the LLM to improve the model’s generalization capabilities when facing unknown domains.A naive training approach involves direct pre-training and SFT using RGB images and constructed QA data.
 However, the high similarity between real faces and PAs in RGB appearance poses significant challenges to this method.
 As shown in Fig.
-4(b), Spoof-Aware Vision Perception (SAVP) combines images preprocessed based on prior knowledge, by extracting predefined local descriptor operators [72], with the original RGB image.
+4(a), Spoof-Aware Vision Perception (SAVP) combines images preprocessed based on prior knowledge, by extracting predefined local descriptor operators (yu2024rethinking), with the original RGB image.
 Our complete model framework is shown in Fig.
-4(c), RGB images and the extracted local descriptor images are fed into the vision encoder to extract vision token $V_{RGB}$ and $V_{SAV}$, respectively.
+4(b), RGB images and the extracted local descriptor images are fed into the vision encoder to extract vision token $V_{RGB}$ and $V_{SAV}$, respectively.
 These features are then processed through the Prompt-Guided Vision Token Masking (PVTM) module, which extracts highly generalizable vision tokens.
@@
 $\mathcal{Y}=\text{MLLM}(V_{align},P)$ (2)
-4.1 Spoof-Aware Vision Perception
+Spoof-Aware Vision Perception
+Bonafide faces and PAs lack distinct discriminative features in RGB-based appearance space, whereas local descriptors (yu2020searching; yu2024rethinking; xie2024fusionmamba) extracted through image preprocessing can enhance their subtle live/spoof clues.
 As shown in Fig.
-4(a), bonafide faces and PAs lack distinct discriminative features in RGB-based appearance space, whereas local descriptors [66, 72] extracted through image preprocessing can enhance their subtle live/spoof clues.
-As shown in Fig.
-4(b), we extract features from the original images using Local Binary Pattern (LBP) [47], Gray, and Histogram of Oriented Gradients (HOG) [11], and concatenate them.
+4(a), we extract features from the original images using Local Binary Pattern (LBP) (pietikainen2010local), Gray, and Histogram of Oriented Gradients (HOG) (dalal2005histograms), and concatenate them.
 LBP and Gray-specific computations are as follows:
 $LBP$ $=\sum_{i=0}^{P-1}s(g_{i}-g_{c})\cdot 2^{i},$ (3)
-$s(x)$ =$=\begin{cases}1&\text{if }x\geq 0,\\ 0&\text{otherwise},\end{cases}$
+[…]
 where $g_{c}$ is the pixel value of the central pixel in the considered neighborhood, and $g_{i}$ represents the pixel values of the $P$ surrounding pixels.
@@
 $V=\text{Concat}[V_{RGB},V_{SAV}])$ (6)
-Figure 5: Prompt-guided Vision Token Masking Module (PVTM)
-4.2 Prompt-Guided Vision Token Masking
+Prompt-Guided Vision Token Masking
 To further enhance the alignment between visual features and text prompts, and alleviate overfitting on spurious correlations, we leverage text prompts to guide token selection after the visual encoder.
 As shown in Fig.
-5, the text tokens extracted from the prompt are pooled [54], and then their similarity with all visual tokens is calculated.
+4(c), the text tokens extracted from the prompt are pooled (song2024moresimpleeffectivetoken), and then their similarity with all visual tokens is calculated.
 We assume that visual tokens with higher similarity are more relevant to subsequent tasks and even for the same image, the important visual tokens may not be consistent across different tasks.
@@
 The remaining tokens are then randomly masked with a probability of $p\%$, reducing the influence of less important tokens while keeping acceptable information loss for the final decision-making process.
-4.3 Training Details
-We employ a two-stage training strategy.
-The first stage is the pre-training stage, where we perform continual pretrain of existing weights.
-In this stage, we align the visual embeddings from a pre-trained vision encoder with the text embeddings using FaceShield-pre10K.
-The second stage is the supervised fine-tuning (SFT) stage, where we apply visual instruction tuning to fully leverage the MLLM’s capabilities across various multimodal tasks within our specific domain using FaceShield-sft45K.
-For LLM tuning, we incorporate LoRA [22] and set the parameters with a hidden dimension of 128 and a scaling factor of 256 in every transformer block.
-In both stages, we employ cross-entropy loss for next-token prediction.
-During the pre-training stage, only the projector after the vision encoder and PVTM are optimized for one epoch.
-In the SFT stage, we fine-tune the LoRA layers in LLM as well as the projector.
-Table 1: Intra-dataset results on coarse-grained classification.
-Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
-Traditional
-ResNet [19] 97.55 2.32
-PatchNet [59] 98.22 1.78
-CoOp [77] 98.73 1.27
-MLLM
-LLaVA [38] 65.54 27.76
-Qwen-VL [4] 51.94 38.70
-Minigpt4 [79] 26.86 65.50
-Bunny [20] 81.20 17.87
-Bunny (fine-tuned) [20] 98.23 1.52
-FaceShield (Ours) 99.41 ± 0.06 0.53 ± 0.06
-5 Experiments and Results
-5.1 Protocols and Evaluation Metrics
-We select 10% of each source dataset in FaceShield-sft45K to form three test subsets (W, S, P).
-For the coarse-grained classification and fine-grained classificationtask, we conduct both intra-dataset and cross-dataset testing.
-For the reasoning, and attack localization tasks, we employ intra-dataset evaluation.
-The intra-dataset test uses all data from the source datasets for training and evaluation on the combined test sets (W&S&P).
-For cross-dataset testing, we select two datasets for pre-training and SFT, and then test using the test set from the remaining dataset.
-For example, we might train on W and S, and test using the test set from P.
-For the coarse-grained classification, fine-grained classification, and reasoning tasks, we use Half Total Error Rate (HTER) [69] or accuracy (ACC) as evaluation metrics.
-Additionally, we evaluate the reasoning process for the reasoning task using BLEU [45], ROUGE-L [34], and METEOR [5], and evaluate the attack localizationtask using AP@40 and AP@50.
-5.2 Implementation Details
-We use Siglip [74] as the visual encoder and Phi-3 [1] as the language foundation model.
+Training Details
+We adopt a two-stage training strategy.
+In the first stage (pretraining), we perform continual pretraining to align visual embeddings from a pretrained vision encoder with text embeddings using FaceShield-pre10K.
+In the second stage (supervised fine-tuning, SFT), we apply visual instruction tuning on FaceShield-sft45K to fully exploit the MLLM’s capabilities across domain-specific multimodal tasks.
+For LLM adaptation, we apply LoRA (hu2021loralowrankadaptationlarge) with a rank of 128 and a scaling factor of 256 for each transformer block.
+Cross-entropy loss is used for next-token prediction in both stages.
+During pretraining, we update only the vision projector and the PVTM module for one epoch.
+In the SFT stage, we fine-tune the LoRA layers in the LLM along with the vision projector.
+Experiments and Results
+Protocols and Evaluation Metrics
+We use 10% of each source dataset in FaceShield-sft45K to construct three test subsets: W, S, and P.
+For coarse-grained classification we perform both intra- and cross-dataset evaluations.
+For fine-grained classification, reasoning and attack localization, we conduct intra-dataset evaluation only.
+In intra-dataset testing, models are trained on all source data and evaluated on the combined test sets (W&S&P).
+For cross-dataset testing, two datasets are used for training (including pretraining and SFT), and the remaining one for testing (e.g., training on W and S, testing on P).For coarse-grained classification, fine-grained classification, and reasoning, we report Half Total Error Rate (HTER) (yu2022deep) and Accuracy (ACC).
+For reasoning, we further evaluate reasoning quality using BLEU (papineni2002bleu), ROUGE-L (lin2004rouge), and METEOR (banerjee2005meteor).
+For attack localization, we report AP@40 and AP@50.
+Implementation Details
+We use Siglip (zhai2023sigmoid) as the visual encoder and Phi-3 (abdin2024phi) as the language foundation model.
 PVTM retains the top 10% of the most important tokens and randomly masks 5% of the tokens in the remaining 90%.
@@
 Each experiment is repeated 10 times on the model, and the final results are reported as the mean ± standard deviation.
+Comparison with Existing Methods
+Coarse-Grained Classification Task
+For the coarse-grained classification task, we compare FaceShield with state-of-the-art FAS methods  (he2016deep; Wang_2022_CVPR; zhou2022learning; zhou2023instance) and open-source MLLMs (liu2023llava; Qwen-VL; zhu2023minigpt; he2024bunny).
+Table 2: Intra-dataset results on coarse-grained classification.
+Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
+Traditional
+ResNet (he2016deep) 97.55 2.32
+PatchNet (Wang_2022_CVPR) 98.22 1.78
+CoOp (zhou2022learning) 98.73 1.27
+MLLM
+LLaVA (liu2023llava) 65.54 27.76
+Qwen-VL (Qwen-VL) 51.94 38.70
+Minigpt4 (zhu2023minigpt) 26.86 65.50
+Bunny (he2024bunny) 81.20 17.87
+Bunny (fine-tuned) (he2024bunny) 98.23 1.52
+FaceShield (Ours) 99.41 ± 0.06 0.53 ± 0.06
 Table 3: Cross-dataset results on coarse-grained classification. $W$, $S$, and $P$ denote WMCA, SiW-Mv2, and PADISI, respectively.
 Method ACC(%) $\uparrow$ HTER(%) $\downarrow$
-W & S P
-ResNet [19] 46.12 50.00
-PatchNet [59] 77.18 22.87
-IADG [78] 72.96 27.01
+W & S → P
+ResNet (he2016deep) 46.12 50.00
+PatchNet (Wang_2022_CVPR) 77.18 22.87
+IADG (zhou2023instance) 72.96 27.01
+FAS-AUG (ijcv24augmentation) 91.7 7.3
 FaceShield (Ours) 93.17 ± 0.22 6.37 ± 0.21
-W & P S
-ResNet [19] 53.36 49.16
-PatchNet [59] 56.16 45.37
-IADG [78] 57.20 42.81
+W & P → S
+ResNet (he2016deep) 53.36 49.16
+PatchNet (Wang_2022_CVPR) 56.16 45.37
+IADG (zhou2023instance) 57.20 42.81
+FAS-AUG (ijcv24augmentation) 88.2 11.7
 FaceShield (Ours) 89.93 ± 0.15 10.3 ± 0.14
-S & P W
-ResNet [19] 74.01 29.75
-PatchNet [59] 78.15 41.50
-IADG [78] 78.55 26.27
+S & P → W
+ResNet (he2016deep) 74.01 29.75
+PatchNet (Wang_2022_CVPR) 78.15 41.50
+IADG (zhou2023instance) 78.55 26.27
+FAS-AUG (ijcv24augmentation) 87.9 13.1
 FaceShield (Ours) 92.56 ± 0.08 5.71 ± 0.08
+Table 4: Cross-dataset result on CASIA-MFSD and Replay-Attack.
+Test Dataset ACC(%) $\uparrow$ HTER(%) $\downarrow$
+S & P & W $\rightarrow$ CASIA-MFSD 90.59 6.37
+S & P & W $\rightarrow$ Replay-Attack 82.42 20.07
 Table 5: Results of fine-grained classification task.
 Method ACC(%) $\uparrow$
-LLaVA [38] 16.39
-Qwen-VL [4] 16.55
-Minigpt4 [79] 19.51
-Bunny [20] 27.03
-Bunny (Fine-tuned) [20] 94.43
+LLaVA (liu2023llava) 16.39
+Qwen-VL (Qwen-VL) 16.55
+Minigpt4 (zhu2023minigpt) 19.51
+Bunny (he2024bunny) 27.03
+Bunny (Fine-tuned) (he2024bunny) 94.43
 FaceShield (Ours) 95.81 ± 0.11
@@
 Method BLEU-1 (%) $\uparrow$ BLEU-2 (%) $\uparrow$ BLEU-3 (%) $\uparrow$ BLEU-4 (%) $\uparrow$ ROUGE-L (%) $\uparrow$ METEOR (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
-LLaVA [38] 45.05 31.75 23.42 17.80 30.51 25.52 37.84 50.11
-Minigpt4 [79] 17.85 7.85 3.53 1.94 27.54 21.88 33.86 50.00
-Qwen-VL [4] 20.92 14.53 11.01 8.77 21.45 12.49 47.64 41.49
-Bunny [20] 33.64 27.12 22.65 19.33 36.74 19.12 50.68 39.73
-Bunny(fine-tuned) [20] 89.57 86.96 84.91 81.29 80.15 51.64 98.56 1.16
+LLaVA (liu2023llava) 45.05 31.75 23.42 17.80 30.51 25.52 37.84 50.11
+Minigpt4 (zhu2023minigpt) 17.85 7.85 3.53 1.94 27.54 21.88 33.86 50.00
+Qwen-VL (Qwen-VL) 20.92 14.53 11.01 8.77 21.45 12.49 47.64 41.49
+Bunny (he2024bunny) 33.64 27.12 22.65 19.33 36.74 19.12 50.68 39.73
+Bunny(fine-tuned) (he2024bunny) 89.57 86.96 84.91 81.29 80.15 51.64 98.56 1.16
 FaceShield (Ours) 90.89 ± 0.14 88.02 ± 0.15 85.75 ± 0.17 83.98 ± 0.19 82.98 ± 0.20 53.10 ± 0.16 99.29 ± 0.04 0.57 ± 0.04
-5.3 Comparison with Existing Methods
-5.3.1 Coarse-Grained Classification Task
-For the coarse-grained classification task, we compare FaceShield with state-of-the-art FAS methods  [19, 59, 77, 78] and open-source MLLM models [38, 4, 79, 20].
+Table 7: Results of the attack localization task with metrics AP@40 and AP@50.
+Method AP@40 (%) $\uparrow$ AP@50 (%) $\uparrow$
+Qwen-VL (Qwen-VL) 2.07 1.49
+Lenna (wei2023lenna) 37.77 35.41
+Sphinx (lin2023sphinxjointmixingweights) 47.86 46.30
+Bunny (he2024bunny) 73.50 71.65
+Bunny(fine-tuned) (he2024bunny) 92.30 89.71
+FaceShield (Ours) 97.78 ± 0.21 95.60 ± 0.19
+Table 8: Ablation study on pretraining w/ or w/o FaceShield-pre10K dataset.
+FaceShield-pre10K Fine-grained Classification Reasoning
+ACC (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
+$\times$ 94.78 $\pm$ 0.19 98.83 $\pm$ 0.06 0.94 $\pm$ 0.05
+$\checkmark$ 95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04
 Intra-dataset Testing.
-Table 1 demonstrates that Our FaceShield significantly outperforms three representative traditional FAS methods [19, 59, 77] in within-domain experiments.
+Table 2 demonstrates that Our FaceShield significantly outperforms three representative traditional FAS methods (he2016deep; Wang_2022_CVPR; zhou2022learning).
 Moreover, our performance greatly exceeds the zero-shot capabilities of general MLLM.
-We also fine-tune the open-source MLLM (i.e., Bunny [20]), selecting RGB images and language data from the dataset to conduct experiments on Bunny.
+We also fine-tune the open-source MLLM (i.e., Bunny (he2024bunny)), selecting RGB images and language data from the dataset to conduct experiments on Bunny.
 We find that FaceShield also surpasses the well-tuned MLLM (Bunny) with 1% HTER decrease.
@@
 FaceShield demonstrates performance far exceeding traditional FAS models in cross-domain scenarios.
-Under the S&PW protocol, it achieves the HTER of 5.72%, showcasing FaceShield’s strong generalization capabilities compared to traditional methods.
-5.3.2 Fine-Grained Classification Task
+Under the S&P→W protocol, it achieves the HTER of 5.72%, showcasing FaceShield’s strong generalization capabilities compared to traditional methods.
+Further results in Table 4 confirm its robustness across unseen datasets like CASIA-MFSD and Replay-Attack.
+Table 9: Ablation results across different tasks.
+SAVP PVTM Coarse-grained classification Fine-grained classification Reasoning Attack localization
+ACC (%) $\uparrow$ HTER (%) $\downarrow$ ACC (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$ AP@40 (%) $\uparrow$ AP@50 (%) $\uparrow$
+$\times$ $\times$ 98.23 1.52 94.43 98.56 1.16 92.30 89.71
+$\times$ $\checkmark$ 98.32 1.83 95.06 98.70 1.04 92.21 90.82
+$\checkmark$ $\times$ 98.73 1.06 94.59 99.41 0.48 97.09 95.21
+$\checkmark$ $\checkmark$ 99.41 $\pm$ 0.06 0.53 $\pm$ 0.06 95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04 97.78 $\pm$ 0.21 95.6 $\pm$ 0.19
+Fine-Grained Classification Task
 Table 5 shows the results under the fine-grained classification task.
-For open-source MLLMs [38, 4, 79, 20], we incorporated 12 types of attacks into the prompt, allowing it to respond with the correct type.
+For open-source MLLMs, we incorporated 12 types of attacks into the prompt, allowing it to respond with the correct type.
 For the fine-tuned MLLM and our FaceShield, we selected keywords from the responses for evaluation.
 It is evident that supervised fine-tuning can significantly improve the model’s performance, with FaceShield achieving the best results.
-Cross-dataset results can be found in the Appendix due to space constraints.
-5.3.3 Reasoning Task
+Reasoning Task
 We also explore the models’ reasoning capacity and Table 6 displays the results for the reasoning task.
@@
 FaceShield not only provides accurate results but also delivers detailed and correct reasoning, effectively enhancing the explainability of FAS methods.
-5.3.4 Attack Localization Task
+Attack Localization Task
 It is vital to locate the spoof regions for explainable FAS, and Table 7 presents the results for the attack localization task.
@@
 It accurately locates attack areas, providing new insights into attack region detection for FAS tasks.
-Table 5: Results of the attack localization task with metrics AP@40 and AP@50.
-Method AP@40 (%) $\uparrow$ AP@50 (%) $\uparrow$
-Qwen-VL [4] 2.07 1.49
-Lenna [63] 37.77 35.41
-Sphinx [35] 47.86 46.30
-Bunny [20] 73.50 71.65
-Bunny(fine-tuned) [20] 92.30 89.71
-FaceShield (Ours) 97.78 ± 0.21 95.60 ± 0.19
-Table 6: Ablation results across different tasks.
-SAVP PVTM Coarse-grained classification Fine-grained classification Reasoning Attack localization
-ACC(%) $\uparrow$ HTER(%) $\downarrow$ ACC(%) $\uparrow$ ACC(%) $\uparrow$ HTER(%) $\downarrow$ AP@40(%) $\uparrow$ AP@50(%) $\uparrow$
-$\times$ $\times$ 98.23 1.52 94.43 98.56 1.16 92.30 89.71
-$\checkmark$ $\times$ 98.73 1.06 94.59 99.41 0.48 97.09 95.21
-$\checkmark$ $\checkmark$ $\mathbf{99.41\pm 0.06}$ $\mathbf{0.53\pm 0.06}$ $\mathbf{95.81\pm 0.11}$ $99.29\pm 0.04$ $0.57\pm 0.04$ $\mathbf{97.78\pm 0.21}$ $\mathbf{95.6\pm 0.19}$
-Figure 6: Comparison of performance after fine-tuning using our proposed dataset on LLaVA [38] and Bunny [20] models.
-5.4 Ablation Study
+Figure 5: Comparison of performance after fine-tuning using our proposed dataset on LLaVA and Bunny models.
+Figure 6: Ablation of visual token masking ratio $p$ in PVTM on three tasks (i.e., coarse- & fine-grained classification, and reasoning)
+Figure 7: Ablation of visual token retain ratio $k$ in PVTM on three tasks (i.e., coarse- & fine-grained classification, and reasoning)
+Ablation Study
 Effectiveness of the proposed datasets.
-We conduct pre-training and SFT with our proposed FaceShield-pre10K and FaceShield-sft45K on LLaVA [38] and Bunny [20] to evaluate the efficacy and generalization of our instruction tuning datasets.
-The results in Fig.
-6 demonstrate that fine-tuning on our dataset significantly enhances performance across three tasks for both LLaVA and Bunny.
-This demonstrates that our dataset can effectively enrich general MLLMs with substantial FAS-related knowledge and boost their capabilities in FAS tasks.
-It also validates the efficacy of our advanced dataset construction pipeline.
-Figure 7: Ablation of visual token masking ratio $p$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
+We conduct pre-training and SFT with our proposed FaceShield-pre10K and FaceShield-sft45K on LLaVA (liu2023llava) and Bunny (he2024bunny) to evaluate the efficacy and generalization of our constructed datasets.
+As shown in Table 8, pretraining with the FaceShield-pre10K dataset significantly improves performance, with fine-grained classification accuracy and reasoning accuracy increasing, while HTER is reduced.
+This demonstrates that pretraining with FaceShield-pre10K enhances the model’s capabilities in FAS-related tasks.
+Additionally, the results in Fig.
+5 show that fine-tuning on our dataset further boosts performance across three tasks for both LLaVA and Bunny.
+This validates the effectiveness of our dataset in enriching MLLMs with FAS-related knowledge and improving their overall performance in FAS tasks, supporting the efficacy of our advanced dataset construction pipeline.
 Effectiveness of SAVP.
@@
 The local live/spoof details within the auxiliary data proves especially valuable for fine-grained attack region detection tasks.
-Figure 8: Ablation of visual token retain ratio $k$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
 Effectiveness of PVTM.
@@
 This might be because masking partial visual tokens may compromise overall image perception and lose information for reasoning.
-We further conduct studies of PVTM with different visual token masking ratios $p$, as shown in Fig.
-7(a) and (b).
-Through our experiments for three (coarse- and fine-grained classification, and attack localization) tasks, we find that masking 5% of the tokens strikes an optimal balance between reducing spurious correlations and preserving essential information.
-However, we cannot find any improvement when masking more tokens due to severe information loss.
-Besides, we also investigate varying proportions of visual token retain ratio $k$, as illustrated in Fig.
-8(a) and (b).
-We preserve 0%, 10%, 20%, and 30% of the tokens, respectively, and then randomly mask $p$=5the remaining tokens.
-The results show that retaining $k$=10visual tokens with strong importance achieves optimal performance.
-However, the more tokens are preserved, the poorer the model performs, suggesting that as the importance of tokens decreases, the likelihood of spurious correlations increases.
-Figure 9: Importance visualization of (a) RGB vision tokens for classification tasks and (b) SAV tokens for attack localization task.
-6 Visualization and Analysis
-Here, we visualize the importance of RGB and SAV tokens.
+We further study PVTM with different visual token masking ratios $p$, as shown in Fig.6.
+Through experiments on three (coarse- and fine-grained classification, and reasoning) tasks, we find that masking 5% of the tokens strikes an optimal balance between reducing spurious correlations and preserving essential information.
+However, no improvement is found when masking more tokens due to severe information loss.
+Additionally, we investigate varying proportions of visual token retain ratio $k$, as illustrated in Fig.7.
+We preserve 0%, 10%, 20%, and 30% of the tokens, respectively, and then randomly mask pp=5% of the remaining tokens.
+The results show that retaining kk=10% visual tokens with strong importance achieves optimal performance.
+However, the more tokens preserved, the poorer the model performs, suggesting that as the importance of tokens decreases, the likelihood of spurious correlations increases.
+Visualization and Analysis
+Figure 8: Importance visualization of SAV tokens for attack localization task.
 Fig.
-9(a) displays the significance of visual tokens in relation to the prompt text for images from three different categories.
-For the coarse- and fine-grained classification tasks, it is evident that the tokens sensitive to prompt are mainly clustered in the facial regions.
-The features of these areas play a decisive role in determining the final outcome.
-Fig.
-9(b) demonstrates how SAV tokens assist in the attack localization task.
-For this task, RGB tokens sometimes fail to provide accurate localization information.
-As shown in the left of Fig.
-9(b), for the eye mask region, the importance visualization of RGB tokens is quite dispersed.
-However, in the SAV token visualization, the significance of the eye attack region is clearly highlighted.
-The right part of Fig.
-9(b) shows how RGB and SAV tokens together provide supportive information for the attack localization task, including the eye mask attack area and the holding hands.
-7 Conclusion
+8 illustrates the visualization obtained after applying SAVP to the attack localization task, highlighting the effectiveness of SAV tokens.
+Compared to RGB tokens, which may suffer from dispersed attention, SAV tokens, once applied, can accurately focus on spoofed regions (e.g., eyeglass areas, hand-held masks).
+These tokens show strong alignment with the ground-truth annotations, producing more focused and interpretable activations around spoof artifacts.
+Additionally, visual tokens in the attack regions exhibit higher similarity scores with the task prompt.
+With PVTM applied, the model is further guided to focus on these deception-relevant areas, emphasizing the critical roles of both SAV tokens and PVTM in improving localization performance.
+Conclusion
 In this paper, we expand the FAS task into four sub-tasks: coarse-grained classification, fine-grained classification, reasoning, and attack localization, and propose the FaceShield, a specialized MLLM tailored for these FAS tasks.
@@
 This paper represents a preliminary exploration of MLLM for FAS task, and future works will focus on incorporating multiple visual modalities and refining the granularity of attack region localization.
+FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with
+Multimodal Large Language Models
+Appendix A Details about FaccShield Dataset
+Data Source
+The FaceShield dataset, comprising FaceShield-pre10K for pretraining and FaceShield-sft45K for supervised fine-tuning (SFT), is constructed using images sourced from three widely recognized datasets: WMCA (8714076), SiW-Mv2 (xiaoguo2022MDFAS), and PADISI (Rostami_2021_ICCV).
+These foundational datasets provide a diverse basis for generating question-answer (QA) pairs that address a variety of tasks in anti-spoofing research.
+In the pretraining phase, the FaceShield-pre10K dataset focuses on generating QA pairs that describe the visual content of the images.
+A total of 9,297 QA pairs were created to establish a foundational understanding of the visual data.
+These pairs enable models to grasp core visual concepts and perform basic reasoning related to spoofing scenarios, forming the groundwork for subsequent learning.
+In contrast, the FaceShield-sft45K dataset, used in the SFT phase, incorporates task-specific annotations to address more complex objectives.
+This dataset contains 45,662 QA pairs, which are categorized into several distinct tasks: coarse-grained classification, fine-grained classification, reasoning, and attack localization.
+These QA pairs are designed to meet practical anti-spoofing requirements, enabling models to handle diverse challenges across various anti-spoofing tasks.
+The FaceShield-pre10K dataset emphasizes descriptive comprehension and fundamental reasoning about image content, while the FaceShield-sft45K dataset advances to specialized, task-driven annotations.
+The distribution of QA pairs is detailed in Table 10 and the category distributions in Fig.
+9.
+Table.
+11 compares face anti-spoofing datasets.
+Unlike traditional datasets with only binary labels for print and replay attacks, FaceShield provides richer annotations (including attack types and bounding boxes) and supports multimodal tasks with large-scale image-caption and QA pairs.
+Figure 9: Comparison of category distributions in FaceShield-pre10K and FaceShield-sft45K datasets.
+Table 10: Dataset Statistics(QA pairs)
+Datasets Source Data Size Total
+WMCA (8714076) 3875
+FaceShield-pre10K SiW-Mv2 (xiaoguo2022MDFAS) 3782 9297
+PADISI (Rostami_2021_ICCV) 1640
+WMCA (8714076) 16776
+FaceShield-sft45K SiW-Mv2 (xiaoguo2022MDFAS) 18096 45662
+PADISI (Rostami_2021_ICCV) 10790
+Table 11: Comparison of existing datasets
+Dataset Attack Types Annotations
+SiW Print, Replay(2) Binary class
+Oulu-NPU Print, Replay(2) Binary class
+CASIA-MFSD Print, Replay(2) Binary class
+Replay-Attack Print, Replay(2) Binary class
+MSU Print, Replay(2) Binary class
+ROSE Print, Replay, PaperMask(3) Binary class
+FaceShield Unified-attack (11 types) QA and bounding boxes
+Data Format
+We constructed the dataset based on the LLAVA framework, as shown in Fig. 10.
+Figure 10: Data format
+Details about FaceShield-pre10K
+Data Generation
+As shown in Fig.
+11, we employed Bunny-Llama-3-8B-V (MLLM) (he2024bunny) to generate high-quality question-answer (QA) pairs aimed at describing the visual content of images.
+The primary objective of this dataset is to equip the MLLM with a foundational understanding of visual attributes relevant to the Face Anti-Spoofing (FAS) task, enabling it to develop basic perceptual capabilities to distinguish between real and spoofed faces.
+Figure 11: FaceShield-pre10Kconstruction pipeline
+The QA generation process is guided by a systematically designed prompt framework comprising three key components.
+First, each image is paired with its corresponding Ground Truth (GT), which denotes the spoofing category (e.g., glasses, mask, makeup).
+The GT serves as contextual information to ensure that the generated QA pairs are semantically aligned with the spoofing type depicted in the image.
+Second, the construction requirements are explicitly defined to standardize the format and content of the QA pairs.
+Each question is concise, typically a single sentence, and designed to elicit descriptions focused on either the environment (e.g., background, lighting) or facial attributes (e.g., gender, facial organ, expression) (eccv25tffas).
+The answer, by contrast, is required to be detailed and comprehensive, integrating observations about both the environment and facial features.
+This approach ensures that the QA pairs are both informative and task-relevant.
+Lastly, the prompts include illustrative examples to provide clear guidance on the desired output structure and level of detail.
+These examples demonstrate the expected question format, the richness required in the answers, and the integration of diverse attributes such as lighting conditions and facial expressions.
+This structured approach ensures that the generated QA pairs are of high quality, semantically relevant, and tailored to the needs of the FAS domain.
+Consequently, the FaceShield-pre10K dataset provides a robust foundation for the MLLM to develop perceptual capabilities, facilitating its adaptation to more complex and specialized tasks in the subsequent fine-tuning phase.
+Data Filtering
+To ensure the quality of the generated QA pairs, we utilized the CLIP (clip) model to evaluate the semantic alignment between the images and their corresponding textual descriptions.
+The filtering process was based on similarity scores computed between each image and its associated text.
+Samples with a similarity score below 15% were considered low-quality and excluded from the dataset, while the rest were retained.
+For text data exceeding the 77-token limit of CLIP (clip), the text was segmented into multiple sentences, and similarity scores were computed for each sentence.
+The final similarity score was calculated as the average of the sentence-level scores.
+For text samples within the token limit, the similarity score was directly computed using the entire text.
+This approach ensured accurate evaluation for longer text samples while maintaining consistency for shorter ones.
+After applying this filtering process, the dataset was reduced from 12091 to 9297 QA pairs.
+This step ensured that the dataset retained only high-quality, semantically consistent data, forming a reliable basis for pretraining tasks in the FAS domain.
+Details about FaceShield-sft45K
+As shown in Fig.
+12, we utilized Bunny-Llama-3-8B-V (he2024bunny) to generate high-quality QA pairs for four tasks: Coarse-grained Classification, Fine-grained Classification, Reasoning, and Attack Localization.
+The prompts for data generation follow a common structure comprising Ground Truth (GT), Requirements, and Examples.
+GT provides contextual information (e.g., real or spoof labels, spoof types, and bounding box coordinates), Requirements specify the desired QA format, and Examples guide generation with representative samples.
+While the structure is consistent, the content is tailored to the objectives of each task.
+Data sizes for each task are detailed in Table 12.
+Figure 12: FaceShield-sft45K construction pipeline
+The construction process, outlined in Algorithm 1, includes generating QA pairs using multimodal input, quality filtering, human review, and data augmentation through rephrasing.
+This pipeline ensures a diverse, high-quality dataset optimized for the specified tasks.
+Table 12: Task and Data Size(QA pairs) for FaceShield-sft45K Dataset
+Dataset Task Data Size
+FaceShield-sft45K Coarse-grained Classification 13076
+Fine-grained Classification 12749
+Reasoning 11082
+Attack Localization 8755
+Multi-task Data generation
+Coarse-grained Classification Task
+The coarse-grained classification task aims to determine whether a face in an image is real or spoofed.
+This task addresses the fundamental binary classification problem central to FAS.
+In this task, the prompts focus on eliciting clear and definitive judgments about the authenticity of the face, supported by reasoning tied to visual features.
+For example, like Fig.
+13, a typical prompt requires the MLLM to generate a question such as, “Is the face in the picture real or spoof?” and an answer like, “The face in the picture is a spoof face.
+This is because the face appears glossy, has an unusual facial reflex, and therefore it is part of an image played on an electronic device.” This task provides the model with foundational knowledge to identify face authenticity.
+Fine-grained Classification Task
+The fine-grained classification task extends the binary classification by requiring the model to identify not only whether the face is real or spoofed but also the specific spoof type (e.g., glasses, fakehead, makeup attack).
+The prompts for this task are designed to ensure the generated QA pairs explicitly address these finer distinctions.
+For example, like Fig.
+14, a question might ask, “What attack type of facial spoof is shown in this image?” with an answer such as, “The image showcases a Glasses Attack, where the subject’s face is cleverly obscured by a pair of glasses.” This task enhances the model’s ability to detect and classify diverse spoofing techniques.
+Reasoning Task
+The reasoning task is designed to enhance the model’s ability to provide structured and interpretable explanations for its judgments, fostering a deeper understanding of FAS scenarios.
+For instance, like Fig.
+15, the question might be phrased as: “Why is the face real or spoofed?” The answer requires the model to systematically evaluate the image based on four key attributes: Facial Lighting, Global Shape Consistency of Facial Features, Sense of Depth and Three-Dimensionality, and Presence of Phone Screen or Paper Edges.
+These attributes are selected because they represent core indicators in FAS tasks.
+Facial Lighting highlights irregularities caused by artificial materials or spoofing media under light.
+Global Shape Consistency assesses whether the facial features align naturally, a key distinction often disrupted in spoofing attempts.
+Sense of Depth and Three-Dimensionality captures the lack of realistic depth commonly found in spoofed faces.
+Presence of Phone Screen or Paper Edges identifies visible edges or surfaces indicative of physical spoofing mediums, such as photographs or displays.
+By methodically analyzing each of these attributes and integrating their observations, the model delivers a comprehensive judgment on the authenticity of the face.
+Attack Localization Task
+The attack localization task focuses on identifying the specific region of a spoofing attack within an image, which is critical for detailed analysis and interpretability in FAS scenarios.
+The prompts require the model to determine whether a face is real or spoofed and, if spoofed, to locate the spoofing region.
+For instance, like Fig.
+16, the generated QA might include a question like, “Can you locate the spoof area of the face in the image?” The corresponding answer should provide the bounding box (bbox) coordinates of the spoofing region (e.g., eyes, mouth) and a concise description of the observed spoofing characteristics.
+This task ensures that the model not only identifies the presence of a spoof but also pinpoints its location, thereby improving interpretability in FAS decision-making.
+Figure 13: Prompt for Coarse-grained Classification Task
+Figure 14: Prompt for Fine-grained Classification Task
+Figure 15: Prompt for Reasoning Task
+Figure 16: Prompt for Attack Localization Task
+Data Filtering
+To ensure the quality and task relevance of the QA pairs in the FaceShield-sft45K dataset, a multi-stage filtering process was employed, incorporating LLM-based evaluation, special words filtering, and human review.
+This pipeline ensured that low-quality or irrelevant QA pairs were systematically identified and removed, resulting in a high-quality dataset suitable for supervised fine-tuning in Face Anti-Spoofing (FAS) tasks.
+LLM-based Filtering
+The first stage of filtering utilizes an LLM to evaluate the semantic consistency between each question and its corresponding answer.
+The LLM was tasked with determining whether the answer directly addressed the question and adhered to the specific requirements of the task.
+QA pairs were flagged for exclusion if:
+  1. 1.
+    The question and answer exhibited semantic misalignment or inconsistency.
+  2. 2.
+    The answer deviated from the task objective or provided irrelevant information.
+This automated evaluation allowed us to efficiently filter out QA pairs that lacked coherence or failed to meet the standards of task relevance.
+Special Words Filtering
+In parallel, a second filtering mechanism was implemented to identify QA pairs containing predefined special words or phrases indicative of low-quality responses.
+These phrases typically signaled either the model’s inability to provide a meaningful answer or responses irrelevant to the task.
+Examples of such phrases include:
+  • “As an AI language model, I cannot…”
+  • “It is difficult to determine…”
+  • “It is not possible to…”
+QA pairs containing such phrases were automatically removed, as they failed to provide informative or useful annotations for the dataset.
+Human Review and Refinement
+Following the automated filtering stages, human reviewers conducted a final review of the retained QA pairs to ensure quality and consistency. This review process focused on:
+  1. 1.
+    Addressing cases where automated filtering may have incorrectly flagged or missed low-quality QA pairs.
+  2. 2.
+    Correcting minor errors, such as grammatical inconsistencies or subtle misalignments between questions and answers.
+  3. 3.
+    Refining edge cases that required nuanced contextual understanding.
+The human review provided an additional layer of quality assurance, ensuring that the dataset met the required standards of accuracy, relevance, and clarity.
+Data Augmentation
+To further increase the diversity and quantity of QA pairs in the FaceShield-sft45K dataset, a systematic data augmentation strategy that leverages the generative capabilities of the LLM was employed.
+As shown in Fig.
+17, this method used the existing QA pairs as input and generated additional semantically equivalent or closely related QA pairs, enhancing linguistic variability while maintaining the original semantic intent.
+The augmentation process was structured as follows: the LLM was provided with a QA pair in a structured format, including both the question and answer.
+The prompt instructed the LLM to act as an expert in data augmentation, generating multiple new QA pairs that conveyed the same meaning as the original but differed in phrasing, word choice, and structural elements.
+The generated outputs were designed to reflect diverse linguistic expressions while ensuring semantic alignment with the original QA pair.
+By applying this augmentation process consistently across the dataset, the FaceShield-sft45K dataset was significantly enriched with variations that introduced natural linguistic diversity.
+This augmentation ensured that the dataset could support robust model training by exposing the models to a wider range of expressions and syntactic patterns, improving their ability to generalize across varied language inputs.
+Figure 17: Prompt for Data augmentation
+Ablation Study
+This section investigates the contributions of key components in the training pipeline, specifically the inclusion of the FaceShield-pre10K dataset during pretraining and the data filtering and augmentation processes during fine-tuning.
+The ablation experiments provide quantitative insights into the significance of these components.
+To evaluate the impact of pretraining on the FaceShield-pre10K dataset, we compared the performance of FaceShield pretrained with and without this dataset.
+As shown in Table 13, pretraining with FaceShield-pre10K improved fine-grained classification accuracy from $94.78\%\pm 0.19$ to $95.81\%\pm 0.11$, and reasoning accuracy from $98.83\%\pm 0.06$ to $99.29\%\pm 0.04$.
+Moreover, the Half Total Error Rate (HTER) was significantly reduced from $0.94\%\pm 0.05$ to $0.57\%\pm 0.04$, highlighting the dataset’s effectiveness in improving both classification and interpretative capabilities.
+The influence of data filtering and augmentation during supervised fine-tuning was analyzed using three configurations, as summarized in Table 14.
+Without any filtering or augmentation, the model achieved a fine-grained classification accuracy of $87.92\%\pm 0.16$, reasoning accuracy of $98.14\%\pm 0.06$, and an HTER of $1.63\%\pm 0.05$.
+Adding data filtering improved these metrics to $90.88\%\pm 0.19$, $98.73\%\pm 0.03$, and $1.24\%\pm 0.08$, respectively.
+When both filtering and augmentation were applied, the model achieved the best performance with fine-grained classification accuracy of $95.81\%\pm 0.11$, reasoning accuracy of $99.29\%\pm 0.04$, and an HTER reduced to $0.57\%\pm 0.04$.
+These results underscore the complementary effects of filtering, which ensures task-relevant and semantically consistent data, and augmentation, which introduces linguistic variability and improves generalization.
+Table 13: Ablation Study on Pretraining w/ or w/o FaceShield-pre10K Dataset.
+FaceShield-pre10K Fine-grained Classification Reasoning
+ACC (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
+$\times$ 94.78 $\pm$ 0.19 98.83 $\pm$ 0.06 0.94 $\pm$ 0.05
+95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04
+Table 14: Ablation Study on Data filtering and Data augmentation.
+Data filtering Data augmentation Fine-grained classification Reasoning
+ACC(%) $\uparrow$ ACC(%) $\uparrow$ HTER(%) $\downarrow$
+$\times$ $\times$ 87.92 $\pm$ 0.16 98.14 $\pm$ 0.06 1.63 $\pm$ 0.05
+$\times$ 90.88 $\pm$ 0.19 98.73 $\pm$ 0.03 1.24 $\pm$ 0.08
+95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04
+Overall, the ablation study demonstrates the critical role of pretraining on FaceShield-pre10K, which provides foundational knowledge for downstream tasks, and the combined use of data filtering and augmentation, which enhances both data quality and linguistic diversity.
+By isolating these components, we confirm the robustness and effectiveness of the proposed training pipeline for multimodal anti-spoofing tasks.
+Appendix B Evaluation Metrics
+To comprehensively evaluate the performance of the proposed model across different tasks, this study employs a range of evaluation metrics, including Half Total Error Rate (HTER) (yu2022deep), Accuracy (ACC), BLEU (papineni2002bleu), ROUGE-L (lin2004rouge), METEOR (banerjee2005meteor), and Average Precision (AP@40 and AP@50).
+The choice of these evaluation metrics is motivated by the specific nature of each task.
+ACC and HTER are commonly used in Face Anti-Spoofing (FAS) tasks, providing a direct measure of model performance in distinguishing between genuine and spoofed faces.
+For the reasoning task, which involves generating long textual content for analysis, metrics such as BLEU, ROUGE-L, and METEOR are suitable as they are widely used in Natural Language Processing (NLP) to evaluate the quality of generated text.
+Lastly, for the localization task, where the goal is to locate the spoof area, we include object detection metrics like AP@40 and AP@50, which are effective in measuring the accuracy of the predicted regions of interest in detecting spoofed areas.
+The detailed definitions and calculations of these metrics are presented as follows:
+Half Total Error Rate (HTER)
+HTER is a critical metric for binary classification tasks, especially in scenarios with imbalanced data. It is defined as:
+$\text{HTER}=\frac{\text{FAR}+\text{FRR}}{2},$ (9)
+where FAR (False Acceptance Rate) represents the proportion of negative samples that are incorrectly classified as positive, and FRR (False Rejection Rate) represents the proportion of positive samples that are incorrectly classified as negative.
+Accuracy (ACC)
+Accuracy evaluates the proportion of correctly classified samples among all samples. It is defined as:
+$\text{ACC}=\frac{\text{TP}+\text{TN}}{\text{Total}},$ (10)
+where TP and TN are the true positives and true negatives, respectively, and Total is the sum of all samples.
+BLEU (Bilingual Evaluation Understudy)
+BLEU is a metric commonly used in natural language generation tasks to evaluate the n-gram overlap between generated and reference texts.
+In this study, we utilize BLEU-1, BLEU-2, BLEU-3, and BLEU-4, which correspond to the 1-gram, 2-gram, 3-gram, and 4-gram match rates, respectively.
+The general formula for BLEU is:
+$\text{BLEU-N}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (11)
+where $N$ is the maximum n-gram length, $p_{n}$ is the proportion of matched n-grams, $w_{n}$ is the weight (usually equally distributed), and BP is the brevity penalty to adjust for short generated texts:
+$\text{BP}=\begin{cases}1&\text{if }c>r,\\ e^{(1-r/c)}&\text{if }c\leq r,\end{cases}$ (12)
+where $c$ and $r$ are the lengths of the generated text and the reference text, respectively.
+ROUGE-L (Recall-Oriented Understudy for Gisting Evaluation)
+ROUGE-L is based on the longest common subsequence (LCS) and measures the similarity between the generated and reference texts in terms of sequential word matches. It is defined as:
+$\text{ROUGE-L}=\frac{\text{LCS}(\text{gen},\text{ref})}{\max(\text{len}(\text{gen}),\text{len}(\text{ref}))},$ (13)
+where LCS is the length of the longest common subsequence between the generated text gen and the reference text ref, and $\text{len}(\cdot)$ denotes the length of the respective text.
+METEOR (Metric for Evaluation of Translation with Explicit ORdering)
+METEOR evaluates the semantic similarity between the generated and reference texts by considering word matching, stemming, and synonymy. It is computed as:
+$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Penalty}),$ (14)
+where $F_{\text{mean}}$ is the harmonic mean of precision and recall, and Penalty is a factor penalizing repeated patterns.
+Appendix C Memory Requirement
+As shown in Tab.
+15, the FaceShield model, with 3.83B parameters and a size of 10.42 GB, allocates 7.14 GB of memory during inference.
+This model achieves a balance between parameter count, model size, and memory allocation, making it suitable for environments with limited resources while maintaining strong performance.
+Model Parameters Model Size Memory Allocated
+FaceShield 3.83B 10.42 GB 7.14 GB
+Table 15: Model Specifications
+Appendix D Detailed Experiments Setting
+Detailed Setting of Model Architectures
+As illustrated in Fig.
+4(a) depicts the baseline model, similar to LLaVA, comprising key components such as the Vision Encoder, Projector, Tokenizer, and LLM.
+This model only accepts RGB modality images as input.
+However, the high visual similarity between genuine faces and presentation attacks (PAs) in the RGB domain presents a significant challenge for this approach.
+In Fig.
+4(b), the input is extended to include auxiliary modality images, which corresponds to the proposed SAVP strategy.
+Specifically, the auxiliary modality images are generated by applying operations such as Gray, HOG, and LBP to the RGB images, thereby enhancing the model’s ability to perceive attack cues.
+In Fig.
+4(c), we introduce the final model, which incorporates the PVTM module.
+This module retains and masks visual tokens based on the similarity between visual tokens and textual tokens, thereby reducing interference and redundancy.
+For a detailed explanation of the PVTM module, refer to Fig.
+4(c).
+Overall, the framework consists of a Siglip (zhai2023sigmoid) encoder for input processing, Phi-3 (abdin2024phi) for inference, and a 2-layer MLP projector (with GELU activation in each layer) for feature mapping and output.
+The SAVP module utilizes dual-stream input to capture richer deception cues, although it may introduce redundancy.
+The PVTM mechanism effectively filters out unnecessary tokens, allowing the model to focus on the most important cues, ultimately improving performance.
+Detailed Setting of Comparison Methods
+In this study, we compare our proposed method with several baseline models, including PatchNet (Wang_2022_CVPR), CoOp (zhou2022learning), IADG (zhou2023instance), LLaVA (liu2023llava), Qwen-VL (Qwen-VL), Minigpt4 (zhu2023minigpt), and Bunny (he2024bunny).
+All of these comparison methods were implemented using the official open-source code, adhering to the specific settings and configurations provided by the respective authors.
+For the finetuned MLLMs, the training process was conducted using the dataset we introduced in this work.
+Appendix E More Examples and of Experimental Results of FaceShield
+To illustrate the effectiveness of the proposed FaceShield model across various tasks, we present example outputs for the coarse-grained classification, fine-grained classification, reasoning, and attack localization tasks in Fig.
+18 to 21.
+These examples highlight the superior performance of FaceShield compared to other models, including LLaVA (liu2023llava), Minigpt4 (zhu2023minigpt), QwenVL (Qwen-VL), and Bunny (he2024bunny).
+Coarse-Grained Classification Task (Fig. 18 and 19)
+In Fig.
+18, FaceShield correctly identifies the face as a spoof by observing critical features such as the unnatural gloss and artificial attributes of the face (e.g., it is a mannequin with painted features).
+Competing models, such as LLaVA and Qwen, provide inconsistent judgments, while MiniGPT-4 fails to determine the authenticity of the face.
+Similarly, in Figure 19, FaceShield recognizes the spoof face as being displayed on a screen, correctly leveraging cues such as unnatural lighting and the screen context.
+Other models, however, misclassify the face as real or fail to reach a conclusion.
+These results demonstrate that FaceShield excels in identifying subtle spoofing indicators, such as lighting anomalies and material inconsistencies, providing reliable binary classification for real and spoofed faces.
+Fine-Grained Classification Task (Fig. 20)
+In Fig.
+20, FaceShield accurately classifies the spoof type as a “Fakehead Attack,” identifying key features such as lifeless eyes and uniform skin tone, which are typical of a mannequin.
+Competing models either fail to identify the correct spoof type or provide incomplete and inconsistent descriptions.
+Reasoning Task (Fig. 22)
+Fig.
+22 demonstrate FaceShield’s ability to provide comprehensive and interpretable reasoning for its decisions.
+In Figure 22, FaceShield uses a structured reasoning framework based on four core attributes: facial lighting, global shape consistency, sense of depth and three-dimensionality, and the presence of edges indicative of spoofing.
+This structured approach allows FaceShield to conclude that the face is a spoof with high confidence.
+Competing models either misclassify the face as real or provide reasoning that lacks depth and critical observations.
+Attack Localization Task (Fig. 21)
+In Fig.
+21, FaceShield precisely locates the spoofing region (e.g., the eyes covered by a paper card) and provides accurate bounding box coordinates.
+In contrast, other models provide vague or incorrect localization results.
+The precise identification of spoofing areas demonstrates FaceShield’s strength in spatial reasoning and interpretability, which are critical for practical anti-spoofing applications.
+Summary of Advantages
+Across all tasks, FaceShield consistently outperforms competing models due to its:
+  • Accurate Identification: Superior coarse-grained and fine-grained classification of spoofing attacks with fewer errors.
+  • Comprehensive Reasoning: Structured and interpretable analysis based on core visual and contextual attributes.
+  • Precise Localization: Accurate and detailed identification of spoofing regions, enabling actionable insights.
+These examples collectively highlight FaceShield’s capability to address the multifaceted challenges in face anti-spoofing with a high degree of accuracy, interpretability, and robustness.
+Furthermore, to highlight FaceShield’s accuracy in classifying on real human face images, we provide additional examples in Fig.
+24, further validating the model’s robust capability in handling genuine face data.
+Additionally, Fig.
+23 presents examples of misclassified cases.
+Analysis shows that frequent errors occur between Flexible and Rigid Masks, as well as Print and Replay attacks.
+These issues stem from insufficient data for Flexible and Rigid Masks, making it difficult to map material properties to attack types, while distinguishing Replay and Print attacks requires more contextual information.
+Figure 18: Example of Coarse-grained Classification Task
+Figure 19: Example of Coarse-grained Classification Task
+Figure 20: Example of Fine-grained Classification Task
+Figure 21: Example of Attack Localization Task
+Figure 22: Example of Reasoning Task
+Figure 23: Example of Misclassification
+Figure 24: Example of Bonafide Face